OpenStudy (anonymous):

The roots of the quadratic equation 2x^2+11x-40=0 are 2e and (f-2).Calculate the possible values of e and f. @ParthKohli

Parth (parthkohli):

Again, you can use Vieta's Formulas if you want to.

Parth (parthkohli):

Or use the quadratic formula.

Parth (parthkohli):

I'd prefer the latter for this one.

OpenStudy (er.mohd.amir):

sum of roots =-11/2 so 2e+f-2=-11/2------------(i) 4e+2f-4=-11 4e+2f=-7 f=(-7-4e)/2 now product of roots=-40/2 2e.(f-2)=-20 put the value from eq (i) of f in this now it become quadratic in f solve for f then put value of f in (i) get value of e.

OpenStudy (rational):

I would solve the roots instead hmm

Parth (parthkohli):

Yes.

Parth (parthkohli):

By the quadratic formula.

OpenStudy (rational):

looks it can be factored by grouping

Parth (parthkohli):

For some reason, it's hard to factorise when you've got multiples of 10. I can only think of the factors that are multiples of 10.

Parth (parthkohli):

I can confirm that my brain is autistic when it comes to factoring. The \(n^4 +4 \), uh...

OpenStudy (rational):

Haha sometimes it requires irritating prime factorization and checking all combinations one by one i think 40*2 = 10*4*2 = 5*2^4

Parth (parthkohli):

That's what I later did. But I can't ever think of it when a number is in a raw form.

OpenStudy (rational):

same wid me too lol i end up with listing all divisors and selecting the right ones by checking each divisor

OpenStudy (rational):

@MARC_ are you following us

OpenStudy (anonymous):

Thank you @ParthKohli @rational @Er.Mohd.AMIR

OpenStudy (rational):

2x^2+11x-40=0 2x^2 + 16x - 5x - 40 = 0 2x(x+8) - 5(x+8) = 0 (x+8)(2x-5) = 0 x = -8, 5/2

OpenStudy (rational):

so the possible values of \(e\) and \(f\) can be obtained by : \(-8 = 2e\) AND \(5/2 = f-2\) or \(-8 = f-2\) AND \(5/2 = 2e\)

OpenStudy (anonymous):

The given answer in the book is \[e=\frac{ 5 }{ 4 }\]and\[f=\frac{ 9 }{ 2 }\]

OpenStudy (rational):

Hmm we're getting : \(\large e = -4,~~\frac{5}{4}\) and \(\large f = -6,~~\frac{9}{2}\)

OpenStudy (anonymous):

oh i forget to type e=-4 and f=-6

OpenStudy (anonymous):

Thnx @rational :)

OpenStudy (rational):

np :) it will be a good exercise to solve this problem below methods and see each method gives the same answer : 1) Vieta's formulas - as started by Er.Mohd.AMIR 2) solving roots first and comparing - we did this 3) do you remember any other ways ?

OpenStudy (anonymous):

can we use completing the square to solve this question @rational ?

OpenStudy (rational):

yes we can solve the roots using below methods 1) factor by grouping 2) completing the square 3) quadratic formula

OpenStudy (rational):

they all fall under second method though ``` 1) Vieta's formulas - as started by Er.Mohd.AMIR 2) solving roots first and comparing - we did this a) factor by grouping b) completing the square c) quadratic formula 3) do you remember any other ways ? ```

OpenStudy (anonymous):

oh ok Thnx @rational :)